OpenStudy (anonymous):

Finding Slope From an Equation y=-5/2x-5

OpenStudy (zale101):



OpenStudy (zale101):

-5/2 is your slope

OpenStudy (anonymous):

thank but how you got -5/2

OpenStudy (zale101):

are you looking for the slope in the equation y=-5/2-5?

OpenStudy (zale101):

y=-5/2x-5 *

OpenStudy (anonymous):

yes

OpenStudy (zale101):

-5/2 replaces m in the standard slope intercept and m is the slope y=mx+b m is the slope

OpenStudy (zale101):

-5 is b which is the y-intercept, where the y value touches the y-axis

OpenStudy (anonymous):

so you y=-x-5 so the slope will be -x

OpenStudy (zale101):

the equation is already written in slope intercept form, that means the slope is already given

OpenStudy (anonymous):

ok thank you

OpenStudy (zale101):

np :) hope i was helpful